40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 19 00:07:16 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 01:11:19 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 14:47:59 2013, manasa, Update, LSC, Green status 
       Reply  Tue Nov 19 17:47:17 2013, Jenne, Update, LSC, PRMI+2arms attempt 
Message ID: 9407     Entry time: Tue Nov 19 01:11:19 2013     In reply to: 9405
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Green status 

I am able to lock the Yarm ALS, but not at the full gain that I should be.  I attribute this to my mediocre alignment of the path on the PSL table.  EDIT: Manasa pointed out that I forgot to set the PSL FSS slow adjust to ~zero, so the PSL temperature was off, so there wasn't really any hope for me last night.

However, I decided that I should write down the ALS locking procedure, as shown to me by Masayuki on 29Oct2013, that is written in one of the Control Room notebooks.  So, here it is.  I will write channel names and DTT template names for the Y arm, but the procedure is the same for both arms.

  1. Lock and align arms using IR.
  2. Lock green beams to arms.
  3. Align green beams to arms.
  4. Check beatnote alignment on PSL table.
  5. Find beatnote by changing end laser temperature (C1:ALS-Y_SLOW_SERVO2_OFFSET) in steps of ~30, watch spectrum analyser for peak.  Easier if arms are locked in IR, but disable LSC system before moving to step 6.  Beatnote should be less than ~50 MHz, and should have a peak height of about -20dBm or more.  When doing 2 arms, be careful that beatnotes of the different arms do not overlap in frequency.  Manasa reminds me that you must also remember to set the PSL FSS SLOW actuator adjust to near zero, to get the PSL back near its nominal temperature.
  6. Check UGF of phase tracker loop.  (DTT template in /users/Templates/ALS/YALS_PT_OLTF.xml) Want UGF to be ~2kHz.  Change C1:ALS0BEATY_FINE_PHASE_GAIN as necessary.
  7. Start the watch script from the ALS screen to watch for lockloss.
  8. Look at the PHASE_OUT spectrum (DTT template in /users/Templates/ALS/ALS.xml).
  9. Clear history of Phase Tracker Loop (clear hist button on C1:ALS-BEATY-FINE_PHASE screen).  Very important to do this before step 10, every time you get to step 10  (i.e. if you lose lock and are starting over)!
  10. Check sign of loop gain by using + or - 0.1 for the gain (C1:ALS-YARM_GAIN).  Beatnote should immediately stop moving if you have the sign right.  Otherwise, it'll zip around (if it does, repeat step 9, step 10).
  11. Turn gain of ALS up to ~15.  Watch the PHASE_OUT spectrum, look for the servo bump.  When you see it, back off the gain a little.  Gain of ~15 is usually about the right ballpark.
  12. Turn on FM 2, 3, 6, 7, 8 of C1:ALS-YARM.  (FM5 should already have been on).
  13. Wait for PHASE_OUT spectrum to come down.  Turn on FM10 of C1:ALS-YARM.
  14. Check UGF of ALS loop (DTT template in /users/Templates/ALS/YALS_OLTF.xml).  Want UGF to be about 150 or 170 Hz (at the peak of the phase bubble).  Adjust C1:ALS-YARM_GAIN as necessary.
  15. ALS is locked!  Use something like the "Scan Arm" script from the ALS screen to find IR resonance, or do whatever measurement you want.  Dataviewer template /users/Templates/Dataviewer_Templates/ALSdtv.xml may be useful.
ELOG V3.1.3-